# Simulations of Interacting Binary Systems - Pathways to Radio Bright GRB Progenitors

### v1

@@ Abstract
Although the association of gamma-ray bursts (GRBs) with massive stellar death is on firm footing, the nature of the progenitor system and the key ingredients required for a massive star to produce a gamma-ray burst remain open questions. Here, we investigate the evolution of a massive star with a closely orbiting compact object companion using the stellar evolution code MESA. In particular, we examine how the companion influences the angular momentum and circumstellar environment near the end of the massive star life. We find that tidal effects can cause the compact object companion to significantly increase the angular momentum of the massive star, for orbital periods in the range of up to days. We model the density profile evolution of the massive star and discuss how tidal interactions may also lead to stripping of the outer stellar envelope in a way that can create an environment around the binary system that deviates from a typical wind density profile. We show how our results depend on the metallicity of the system, initial spin of the star, mass ratio, as well as accretion and dynamo prescriptions in the simulations. We conclude that these systems may be viable progenitors for radio-bright, long gamma-ray bursts.
@@ 1 Introduction
In this paper, we test the hypothesis that long duration, radio-bright GRBs could result from a massive star collapsing in an interacting binary system with a closely orbiting compact object companion. In particular, we explore the influence of a black hole companion over a range of phase space varying orbital period, metallicity, initial spin of the star, mass ratio, accretion rate, magnetic fields, and numerical stopping conditions. We use the well-developed, open-source stellar evolution code Modules for Experiments in Stellar Astrophysics (MESA) to perform this study. We present the parameter space in which this system will produce a highly-spinning massive star with a potentially dense extended circumburst environment - exactly the conditions needed to produce a radio-bright GRB.
Our paper is organized as follows: In §2, we describe the code and simulation set-up (with a convergence study discussed in our Appendix). In §3, we present our results and show the range of initial conditions that lead to a substantial spin up of the massive star in our systems. Additionally, we provide representative density profiles of the massive star and discuss the implications for the density profile of the circumbinary medium. In §4, we present discussion and caveats to our results. In §5, we summarize our main conclusions and discuss future work to support this effort.
@@ 2 Code and Simulation Set-up
We use the one-dimensional stellar evolution code Modules for Experiments in Stellar Astrophysics (MESA) for our simulations (Paxton et al., 2011, 2013, 2015, 2018, 2019; Paxton, 2021; Jermyn et al., 2023)111We use MESA release r21.12.1.. MESA solves the coupled structure and composition equations simultaneously. The MESA equation of state (EOS) is a blend of the OPAL (Rogers & Nayfonov, 2002), SCVH (Saumon et al., 1995), FreeEOS (Irwin, 2004), HELM (Timmes & Swesty, 2000), PC (Potekhin & Chabrier, 2010), and Skye (Jermyn et al., 2021b) EOSes. Radiative opacities are primarily from OPAL (Iglesias & Rogers, 1993, 1996), with low-temperature data from Ferguson et al. (2005) and the high-temperature, and the Compton-scattering dominated regime by Poutanen (2017). Electron conduction opacities are from Cassisi et al. (2007) and Blouin et al. (2020). Nuclear reaction rates are from JINA REACLIB (Cyburt et al., 2010), NACRE (Angulo et al., 1999) and additional tabulated weak reaction rates Fuller et al. (1985); Oda et al. (1994); Langanke & Martínez-Pinedo (2000). Screening is included via the prescription of Chugunov et al. (2007). Thermal neutrino loss rates are from Itoh et al. (1996). Roche lobe radii in binary systems are computed using the fit of Eggleton (1983). Mass transfer rates in Roche lobe overflowing binary systems are determined following the prescription of Ritter (1988).
We employ the binary module in MESA because it enables simultaneous evolution of an interacting pair of stars undergoing transfer of mass and angular momentum (Paxton et al., 2015). We use the stellar rotation and tidal interaction treatment for the binary evolution from Marchant (2019) and Mink (2019). With these tools, we investigate the tidal regime where a black hole (BH) companion significantly influences the evolution of a massive star prior to collapse. Again, we are in particular searching for the parameter space for which the companion can spin up the massive star, and also produce tidal effects that may lead to a denser (relative to a single massive star), extended circumbinary medium into which the GRB jet will travel. We describe the details of our simulation set-up below.
Model name | [days] | [g cm-2 s-1] | [g cm-2 s-1] | ||||
MS25_BH15_T1.5_Z1.0e-4 | 25 | 15 | 1.5 | 18.86 | |||
MS25_BH15_T2.0_Z1.0e-4 | 25 | 15 | 2.0 | 22.85 | |||
MS25_BH15_T2.5_Z1.0e-4 | 25 | 15 | 2.5 | 26.51 | |||
MS25_BH15_T3.0_Z1.0e-4 | 25 | 15 | 3.0 | 29.94 | |||
MS25_BH15_T4.0_Z1.0e-4 | 25 | 15 | 4.0 | 36.27 | |||
MS25_BH15_T1.5_Z1.0e-2 | 25 | 15 | 1.5 | 18.86 | |||
MS25_BH15_T2.0_Z1.0e-2 | 25 | 15 | 2.0 | 22.85 | |||
MS25_BH15_T2.5_Z1.0e-2 | 25 | 15 | 2.5 | 26.51 | |||
MS25_BH15_T3.0_Z1.0e-2 | 25 | 15 | 3.0 | 29.94 | |||
MS25_BH15_T4.0_Z1.0e-2 | 25 | 15 | 4.0 | 36.27 | |||
MS25_BH10_T1.5_Z1.0e-4 | 25 | 10 | 1.5 | 18.04 | |||
MS25_BH10_T3.0_Z1.0e-4 | 25 | 10 | 3.0 | 28.63 | |||
MS25_BH10_T1.5_Z1.0e-2 | 25 | 10 | 1.5 | 18.04 | |||
MS25_BH10_T3.0_Z1.0e-2 | 25 | 10 | 3.0 | 28.63 |
@@ 2.1 MESA Input Physics
We evolve a Zero Aged Main Sequence (ZAMS) star with a point mass companion at different mass ratios, orbital periods222We consider orbital periods of the BH that lead to significant tidal interaction but do not consider common envelope or merger scenarios., metallicities, stellar rotations, and mass/angular momentum transfer schemes. Table 1 gives the choice of parameters for our simulations.
@@
Mass ratio - We consider a 25 massive star with either a 10 or 15 BH companion as physically viable systems, created in situ or by dynamical capture. Systems with these mass ratios have been observed in X-ray binaries (Kelley et al., 1983; Levine et al., 1993, 2000; Tauris & van den Heuvel, 2006; Falanga et al., 2015). In future work, we perform population synthesis models (see, e.g., Cason et al. 2024) to place constraints on the numbers of these systems we expect in our universe. However, for now, we assume these systems exist in sufficient numbers and focus on their individual evolution.
Orbital Period - We consider orbital periods (and subsequently orbital separations) where the BH companion may tidally influence the massive star at a distance beyond the star’s radius outside of the Roche lobe radius of the massive star. Furthermore, at these separations, a Newtonian treatment to the gravitational interaction of the binary is sufficient because the general relativistic corrections to the orbit and tide are very small.
@@
Spin Synchronization - We use the stellar rotation (in binary systems) settings in Marchant (2019), who model systems similar to what we are considering here. In particular, we turn on a spin synchronization flag which enables the tidal spin up of a massive star by its companion. We keep the default of no initial relaxation of the stellar rotation to the orbital period.
@@
Massive star initial rotation - We consider a range of zero and non-zero initial rotations for the massive star. We initialize massive star non-zero rotation at a velocity given by the binary evolution equations in Hurley et al. (2000) and Hurley et al. (2002), used in the binary population synthesis code COSMIC Breivik et al. (2020). We take specific values from the COSMIC output discussed in Kenoly et al. (2023). Additionally, we consider higher velocity initial conditions for the massive star at half of the break-up velocity.
@@
Mass transfer - For the simulations in this paper, the BH companion is primarily beyond the Roche lobe of the massive star and outside the regime where the choice of mass transfer scheme between the massive star and its companion is important. We employ the default setting for MESA (a Kolb-Ritter (Kolb & Ritter, 1990) mass transfer scheme) and further constrain the mass transfer to be lower than the Eddington limit. We note that there are many subtleties that come into play when choosing the mass transfer prescriptions. For example, Cehula & Pejcha (2023) find an intermediate solution between optically thin and optically thick regimes for Roche lobe overflow; their model finds a mass transfer rate about two times smaller than Kolb & Ritter (1990) and four times smaller than Marchant (2019) for a 30 , low metallicity star. However, again, because we are considering encounters where the initial separation distances are larger than the Roche lobe radius, these subtleties generally do not influence our main results.
We set the prescription for angular momentum loss through mass transfer according to the results of de Mink et al. (2013); Mink (2019); Marchant (2019), who carefully examine how angular momentum is lost and/or transferred between stars in a binary system, for different evolutionary stages, chemical composition, masses, separations, and magnetic field configurations.
@@
Spruit-Taylor Dynamo - Through magnetic coupling between layers in the stellar interior, the Spruit-Taylor dynamo (Spruit, 2002) is a possible pathway to transport angular momentum within the massive star, potentially leading to a spin down. However, as we discuss below, its effect is more pronounced in stars less massive than those we consider here. Nonetheless, we have run simulations with this flag both on and off and find a negligible impact on our results (we discuss this further below).
@@
Rotation in Convection Zone - In the case of rotation in the convection zone, angular momentum can be be transported through various processes internal to the star (Augustson & Mathis, 2019). These processes are controlled by the diffusion coefficients of Goldreich-Schubert-Fricke, Eddington-Sweet circulation, secular shear instability, and dynamical shear instability (see, e.g. Heger et al., 2000, for a description). We use the default MESA settings for rotation in the convection zone as well as the diffusion coefficients.
@@
Stopping Condition - We identify the end point for our simulations as the stage prior to collapse when the carbon-oxygen core forms at 6-8 Myr with a corresponding rapid increase in central temperature and density within the star. We report the maximum angular momentum transfer to the star prior to this stage in the simulated evolution.
@@
In the following, we present our set of binary parameters that lead to viable progenitors of radio bright lGRBs, particularly through significant stellar spin up over the massive star lifetime due to interaction with its companion. In our numerical simulations, we track the evolution of quantities that describe the dynamics of the binary system such as stellar and BH mass, orbital separation of the binary, and change in spin, orbital, and total angular momentum as well as stellar quantities through radial profiles in density and temperature. In the Appendix, we discuss convergence studies of representative models MS25_BH15_T3.0_Z1.0e-4 and MS25_BH15_T3.0_Z1.0e-2 of the tidal interaction. We show that these quantities do not change appreciably over stellar lifetime and model number at increasing mass resolution. Additionally, we identify the end point in our simulations prior to collapse in tracking the evolution of the spin up of the massive star with the formation of the carbon-oxygen core.
@@ 3 Results
3.1 Spin up of the Massive Star
For the binary parameters given in Table 1, our numerical simulations show a significant spin up of the massive star due to the tidal interaction with a BH, consistent with analytic estimates for long GRB binary progenitors (Lloyd-Ronning, 2022). In the following, we describe how these changes in spin angular momentum for the massive star depend on stellar metallicity, massive star to BH mass ratio, and initial stellar spin.
3.1.1 Dependence on Metallicity
In Fig. 1, we show the change in angular momentum of the massive star with initial orbital period and corresponding orbital separation for two stellar metallicities at (left panels) and (right panels). In the top row, we give the change in spin angular momentum or extent to which the star is spun up over the lifetime of the star, in units of initial total angular momentum . These interactions increase in spin up for shorter orbital periods, . This is consistent with a characterization of the tidal encounter through a strength parameter (Press & Teukolsky, 1977), for circular orbits at stellar-BH separation distance ,
for orbital timescale and stellar timescale . For a given star, encounters with smaller orbital periods will give rise to smaller or stronger tidal interactions. For our simulations in Table 1,
we consider binaries in the weak or partial disruption regime . Note that the limit corresponds to the tidal disruption limit where the dynamical time scales for the orbit and star are comparable. Our tidal calculations relating the spin up of massive star to orbital period for using the detailed stellar model described in Sec. 2,
agree with results of the spin up of polytropic stars in the weak tidal encounter regime for (Kochanek, 1992; Khokhlov et al., 1993; Cheng & Evans, 2013). Importantly, we find that for a given mass and initial orbital separation , massive star with lower metallicity (left panel) give rise to lower spin up than higher metallicity (right panel). This decrease in spin is consistent with the encounter strength parameter in Eq. 1, where stars with lower metallicity are more compact and give rise to weaker encounters (larger ) at fixed and orbital separation . In the bottom row, we show the evolution of orbital separation in units of . For all of these encounters, the massive star is tidally captured, with a decrease in . This is consistent with the orbital separation to tidal radius and mass ratio regime described in the parameter study simulating the interaction between stellar BH and main-sequence stars in Kremer et al. (2022).

### v2

@@ Abstract
Although the association of gamma-ray bursts with massive stellar death is on firm footing, the nature of the progenitor system and the key ingredients required for a massive star to produce a gamma-ray burst remain open questions. Here, we investigate the evolution of a massive star with a closely orbiting black hole using the stellar evolution code MESA. In particular, we examine how the companion influences the angular momentum of the massive star as it evolves over its lifetime under a range of orbital periods where the tidal effects are significant. We consider dependence on stellar metallicity, mass ratio, initial stellar spin, as well as accretion and dynamo prescriptions while tracking the mass loss and angular momentum in the simulations. We conclude that massive star-black hole binary systems with low metallicity massive stars under an initial orbital period range of 20-50 days may be viable GRB progenitors with post-collapse black hole masses of 5-10 and a black hole spin parameter .
@@ 1 Introduction
In this paper, we examine the influence of a black hole (BH) companion on a massive star over a range of phase space varying orbital period, metallicity, initial spin of the star, and mass ratio and discuss the implications this has for the final angular momentum of the star and the potential GRB it may produce. We identify a particular region of the binary parameter space where the massive star is subject to negligible mass loss from both wind and Roche-lobe overflow (RLOF) as the tidal effects significantly increase the spin angular momentum prior to stellar collapse at the end of its evolution.
We use the well-developed, open-source stellar evolution code Modules for Experiments in Stellar Astrophysics (MESA) to perform this study. We present the parameter space in which this system will produce a highly-spinning massive star.
Previous work in stellar evolution modeling of these binary progenitors has provided many important details on whether or not the tidal interaction is significant enough to spin up a star under a wide range of parameter space. In a thorough investigation of mass and angular momentum loss, Detmers et al. (2008) find that the best candidates for Wolf-Rayet stars with a compact object companion are at low stellar metallicity, where stellar winds are weak and the orbital separations are stable without merger. Marchant et al. (2016) and Mandel & de Mink (2016) consider the interaction of massive stars in tight binaries under chemically homogenous evolution merger scenarios characterized by mixing and wind-driven mass loss due to rotation and tides. In addition there have been detailed studies looking at the influence of a low mass star on a massive star in a tight binary, exploring the conditions underwhich the low mass star can spin up the resultant black hole that the massive star produces upon collapse, but without merging (Moreno Méndez, 2022). In this work, we investigate a new region of parameter space in our choice of masses, stellar metallicities, and wide separations with orbital periods up to 100 days while exploring the limit of tidal interactions without significant mass loss due to RLOF.
Our paper is organized as follows: In §2, we describe the code and simulation set-up (with a convergence study discussed in our Appendix). In §3, we present our results and show the range of initial conditions that lead to a substantial spin up of the massive star in our systems. Additionally, we provide representative density profiles of the massive star and briefly discuss the implications for the density profile of the circumbinary medium. In §4, we present discussion and caveats to our results. In §5, we summarize our main conclusions and discuss future work to support this effort.
@@ 2 Code and Simulation Set-up
We use the one-dimensional stellar evolution code Modules for Experiments in Stellar Astrophysics (MESA) for our simulations (Paxton et al., 2011, 2013, 2015, 2018, 2019; Paxton, 2021; Jermyn et al., 2023)111We use MESA release r21.12.1.. MESA solves the coupled structure and composition equations simultaneously. The MESA equation of state (EOS) is a blend of the OPAL (Rogers & Nayfonov, 2002), SCVH (Saumon et al., 1995), FreeEOS (Irwin, 2004), HELM (Timmes & Swesty, 2000), PC (Potekhin & Chabrier, 2010), and Skye (Jermyn et al., 2021b) EOSes. Radiative opacities are primarily from OPAL (Iglesias & Rogers, 1993, 1996), with low-temperature data from Ferguson et al. (2005) and the high-temperature, and the Compton-scattering dominated regime by Poutanen (2017). Electron conduction opacities are from Cassisi et al. (2007) and Blouin et al. (2020). Nuclear reaction rates are from JINA REACLIB (Cyburt et al., 2010), NACRE (Angulo et al., 1999) and additional tabulated weak reaction rates Fuller et al. (1985); Oda et al. (1994); Langanke & Martínez-Pinedo (2000). Screening is included via the prescription of Chugunov et al. (2007). Thermal neutrino loss rates are from Itoh et al. (1996). Roche lobe radii in binary systems are computed using the fit of Eggleton (1983). Mass transfer rates in Roche lobe overflowing binary systems are determined following the prescription of Ritter (1988). We employ the binary module in MESA because it enables simultaneous evolution of an interacting pair of stars undergoing transfer of mass and angular momentum (Paxton et al., 2015). We use the stellar rotation and tidal interaction treatment for the binary evolution from Marchant (2019) and Mink (2019). With these tools, we investigate the tidal regime where a BH companion significantly influences the evolution of a massive star prior to collapse. We describe the details of our simulation set-up below.
| [days] | ||||
| 15.0 | 4.445 | 2.0 | 2.285 | |
| 15.0 | 4.445 | 3.0 | 2.994 | |
| 15.0 | 4.445 | 10.0 | 6.680 | |
| 15.0 | 4.445 | 15.0 | 8.754 | |
| 15.0 | 4.445 | 20.0 | 1.060 | |
| 15.0 | 4.445 | 21.0 | 1.095 | |
| 15.0 | 4.445 | 22.0 | 1.130 | |
| 15.0 | 4.445 | 23.0 | 1.164 | |
| 15.0 | 4.445 | 24.0 | 1.197 | |
| 15.0 | 4.445 | 25.0 | 1.231 | |
| 15.0 | 4.445 | 30.0 | 1.390 | |
| 15.0 | 4.445 | 40.0 | 1.683 | |
| 15.0 | 4.445 | 50.0 | 1.953 | |
| 15.0 | 4.445 | 60.0 | 2.206 | |
| 15.0 | 4.445 | 100.0 | 3.101 | |
| 15.0 | 6.058 | 2.0 | 2.285 | |
| 15.0 | 6.058 | 3.0 | 2.994 | |
| 15.0 | 6.058 | 10.0 | 6.680 | |
| 15.0 | 6.058 | 15.0 | 8.754 | |
| 15.0 | 6.058 | 20.0 | 1.060 | |
| 15.0 | 6.058 | 21.0 | 1.095 | |
| 15.0 | 6.058 | 22.0 | 1.130 | |
| 15.0 | 6.058 | 23.0 | 1.164 | |
| 15.0 | 6.058 | 24.0 | 1.197 | |
| 15.0 | 6.058 | 25.0 | 1.231 | |
| 15.0 | 6.058 | 30.0 | 1.390 | |
| 15.0 | 6.058 | 40.0 | 1.683 | |
| 15.0 | 6.058 | 50.0 | 1.953 | |
| 15.0 | 6.058 | 60.0 | 2.206 | |
| 15.0 | 6.058 | 100.0 | 3.101 | |
| 10.0 | 4.445 | 10.0 | 6.389 | |
| 10.0 | 4.445 | 20.0 | 1.014 | |
| 10.0 | 4.445 | 30.0 | 1.329 | |
| 10.0 | 4.445 | 40.0 | 1.610 | |
| 10.0 | 4.445 | 50.0 | 1.868 | |
| 10.0 | 4.445 | 60.0 | 2.110 | |
| 10.0 | 6.058 | 10.0 | 6.389 | |
| 10.0 | 6.058 | 20.0 | 1.014 | |
| 10.0 | 6.058 | 30.0 | 1.329 | |
| 10.0 | 6.058 | 40.0 | 1.610 | |
| 10.0 | 6.058 | 50.0 | 1.868 | |
| 10.0 | 6.058 | 60.0 | 2.110 |
@@ 2.1 MESA Input Physics
We evolve a Zero Aged Main Sequence (ZAMS) star in a circular orbit with a point mass BH at different mass ratios, orbital periods222We consider orbital periods of the BH that lead to significant tidal interaction but do not consider common envelope or merger scenarios., metallicities, stellar rotations, and mass/angular momentum transfer schemes. Table 1 gives the choice of parameters for our simulations.
@@
Mass ratio - We consider a 25 massive star with either a 10 or 15 BH companion as physically viable systems, created in situ or by dynamical capture. Systems like these, with these particular mass ratios, have been observed in X-ray binaries (Kelley et al., 1983; Levine et al., 1993, 2000; Tauris & van den Heuvel, 2006; Falanga et al., 2015). In future work, we perform population synthesis models (A. Cason et al. in prep) to place constraints on the numbers of these systems we expect in our universe. However, for now, we assume these systems exist in sufficient numbers and focus on their individual evolution.
Orbital Period - We investigate orbital periods (and subsequently orbital separations) where the BH companion may tidally influence the massive star at a distance beyond the star’s radius outside of the Roche lobe radius of the massive star. Furthermore, at these separations, a Newtonian treatment to the gravitational interaction of the binary is sufficient because the general relativistic corrections to the orbit and tide are very small.
@@
Spin Synchronization - The binary module in MESA assumes that the star’s rotational axis is perpendicular to the orbital plane, enabling tidal interactions to contribute to the star’s spin-up through angular momentum transfer. MESA uses a modified scheme to solve for the star’s angular frequency as a function of time, following the formalism of Hut (1981) for a point-like mass companion while accounting for differential rotating stars (Paxton et al., 2015; Marchant, 2019). We do not initially relax the stellar rotation to the orbital period.
@@
Massive star initial rotation - We consider a range of stable, initial rotations for the massive star under the limit of stellar break-up velocity. We initialize massive star non-zero rotation at a velocity given by the binary evolution equations in Hurley et al. (2000) and Hurley et al. (2002a), used in the binary population synthesis code COSMIC Breivik et al. (2020). We initialize our simulations with zero velocity as a standard point in investigating the effects of spin. We also use two additional criteria for the initial velocity: The first ( km/s ) is set by the binary stellar evolution equations of Hurley et al. (2002b) and studied in Kenoly et al. (2023), who used the COSMIC population synthesis code (Breivik et al., 2020). The other is half of the critical velocity , defined as the velocity at which the star’s rotation rate will cause break up.
@@
Mass transfer - For the simulations in this paper, the BH companion is initially beyond the Roche lobe of the massive star and outside the regime where the choice of mass transfer scheme between the massive star and its companion is important. We employ a Kolb-Ritter (Kolb & Ritter, 1990; Paxton et al., 2015) mass transfer scheme, which accounts for stellar parameters such as finite atmosphere scale height and detailed structure of the outer layers. We further constrain the mass transfer to be lower than the Eddington limit. We note that there are many subtleties that come into play when choosing the mass transfer prescriptions. For example, Cehula & Pejcha (2023) find an intermediate solution between optically thin and optically thick regimes for RLOF; their model finds a mass transfer rate about two times smaller than Kolb & Ritter (1990) and four times smaller than Marchant et al. (2016) for a 30 , low metallicity star. However, again, because we are considering encounters where the initial separation distances are larger than the Roche lobe radius, these subtleties generally do not influence our main results.
Furthermore, we note that in our simulations non-conservative mass transfer occurs such that mass from the massive star due to winds or RLOF does not necessarily end up accreted onto the black hole. In MESA, the corresponding loss in angular momentum is calculated from the net mass loss (Paxton et al., 2015; Marchant, 2019; Mink, 2019).
@@
Spruit-Tayler Dynamo - Through magnetic coupling between layers in the stellar interior, the Spruit-Tayler dynamo (Spruit, 2002; Fuller et al., 2019; Ma & Fuller, 2019) is a possible pathway to transport angular momentum within the massive star, potentially leading to a spin down. However, as we discuss below, its effect is more pronounced in stars less massive than those we consider here (see also the analytic arguments and discussion in Moreno Méndez, 2014, who show that for slowly spinning massive stars the slow-down from a Spruit-Taylor dynamo may be avoided). Nonetheless, we have run simulations with this flag both on and off and find a negligible impact on our results (we discuss this further below).
@@
Rotation in Convection Zone - In the case of rotation in the convection zone, angular momentum can be be transported through various processes internal to the star (Augustson & Mathis, 2019). These processes are controlled by the diffusion coefficients of Goldreich-Schubert-Fricke (GSF), Eddington-Sweet (ES) circulation, secular shear instability (SSI), and dynamical shear instability(DSI). The GSF (Goldreich & Schubert, 1967; Fricke, 1968) coefficient quantifies how angular momentum is transported in the interiors of stars. ES occurs when there are imbalances of pressure and temperature inside the interior of the star. SSI is a long-term instability in rotating stars where thermal diffusion weakens stratification, allowing shear-driven turbulence to redistribute angular momentum and mix elements. DSI occurs in rotating stars when shear forces exceed stabilizing buoyancy, triggering turbulence and rapid mixing on a dynamical timescale (Heger et al., 2000).
@@
Termination condition - We identify the end point for our simulations as the stage prior to collapse when the carbon-oxygen core forms at 6-8 Myr with a corresponding rapid increase in central temperature and density within the star. We also track the evolution of the helium core and central mass ratio of helium-4 and note that our stopping criteria for the simulations is consistent with this end point. We report the final angular momentum transfer to the star at this stage in the simulated evolution.
@@
In the following, we present our set of binary parameters that could lead to viable progenitors of radio bright lGRBs, particularly through significant stellar spin up over the massive star lifetime due to interaction with its companion. In our numerical simulations, we track the evolution of quantities that describe the dynamics of the binary system such as stellar and BH mass, orbital separation of the binary, and change in spin, orbital, and total angular momentum as well as stellar quantities such as mass loss due to wind and RLOF, stellar radius, RLOF radius, and rotation. We report the spin angular momentum and total mass loss for massive stars at the final stage of evolution indicated by our stopping criteria. In the Appendix, we discuss convergence studies of representative models of the binary interaction between a BH and massive star at an initial orbital period of three days to test our schemes where significant mass loss occurs. We show that our quantities of interest do not change appreciably over stellar lifetime and model number at increasing mass resolution.
@@ 3 Results
For the binary parameters given in Tab. 1, our numerical simulations show a particular region with significant spin up of the massive star due to the tidal interaction with a BH, consistent with analytic estimates for long GRB binary progenitors (Lloyd-Ronning, 2022). In Fig. 1, we show the dependence of final stellar spin angular momentum and mass loss on initial orbital period with models given in Tab. 1 for initially non-rotating stars. In Fig. 2 and Fig. 3, we provide an analysis of these models at initially three day and twenty-one day orbital periods and show the evolution of stellar radius, RLOF radius, mass loss due to stellar wind and RLOF, spin and orbital angular momentum, and our termination conditions in tracking the formation of the helium and carbon-oxygen core mass and central helium-4 mass fraction. In the following, we describe our results in detail with regards to the identification of the parameter space of viable progenitor binary systems, mass loss in late stage evolution due to RLOF, angular momentum, termination conditions, metallicity, mass-ratio, rotation, and additional tests and caveats.
3.1 Viable progenitor binary systems with significant stellar spin-up and negligible mass loss
We identify a region of parameter space where the tidal interaction in the binary leads to significant stellar spin-up such that after the massive star collapses, its remnant BH-disk system may retain enough angular momentum to launch a jet via Blandford-Znajek process (Blandford & Znajek, 1977). For a viable GRB from such a progenitor binary system, we estimate the lower limit for the final spin angular momentum as gcm-2s-1 using a BH mass of spin of 0.5. For a BH mass of , the lower limit is smaller by a factor of 4. In the first row of Fig. 1, we show a region of parameter space in initial orbital period , stellar to BH mass ratios, and stellar metallicity where the final stellar spin-up is above the lower limit of a viable GRB progenitor with BH mass of (solid line) and (dotted line). We identify 25, low metallicity () massive stars interacting with both 15 and 10 BH for an initial orbital period of 20-50 days as viable progenitor binary systems. For this particular region of binary parameter space, we note negligible final total mass loss on the order of a factor of of the initial mass of the star. Outside of this region, non-viable binaries are characterized by significant final mass loss close to a factor of unity or at an orbital separation where the tidal interaction is negligible ( 60). Additionally, we note very slight decreases in orbital separation at the level of for simulations with negligible mass loss including the viable progenitor binary systems. This is consistent with Detmers et al. (2008), where it is shown that, while the spin-up process itself may work, significant mass loss due to winds and RLOF negate the spin-up effect through the corresponding loss in stellar angular momentum. Additionally, our results are comparable to the orbital separation to tidal radius and mass ratio regime described in the parameter study simulating the interaction between stellar BH and main-sequence stars in Kremer et al. (2022).
3.2 Roche-lobe overflow in late stage stellar evolution
For all simulations in Tab. 1, the radius of the massive star increases significantly in the late-stages of stellar evolution until termination at Myr. While the RLOF radius remains relatively constant, there is an increase in the ratio between stellar to RLOF radius, leading to significant mass loss. We show that the primary drivers of mass loss are due to stellar winds and RLOF in simulations between a massive star and a 15 BH with an initial orbital period of three days in Fig. 2 and twenty-one days in Fig. 3. In the first row of both figures, we compare the ratio of stellar to RLOF radius (solid line), stellar to initial stellar radius (dashed line), and Roche flow overflow to initial Roche flow overflow radius (dashed dotted line). We note that for the majority of the evolution the ratio is larger for a three-day orbital period than a twenty-one day orbital period because of the initial orbital separation of the former is smaller than the latter in Tab. 1.
For simulations with an initial three day orbital period in Fig. 2, the time of increase in stellar radius close to termination is coincident with a sharp increase in total mass loss , given in the second row. Furthermore, we compare the evolution of total mass loss , mass loss due to stellar wind , and mass loss due to RLOF in units of initial stellar mass . Before the sharp increase in stellar radius, less than Myr, the mass loss is driven by stellar winds. Afterwards, the mass loss is dominated by RLOF. This is the case for the range of stellar metallicities given in both columns.
However, for simulations with an initial twenty-one day orbital period in Fig. 2, we find different behavior at low and high metallicity. In comparing both columns of the second row, we see that at higher metallicity () RLOF is the main driver of mass loss whereas for lower metallicity () its contribution is less than the mass loss due to stellar wind. We note dips in the late-stage evolution of the expansion in stellar radius for simulations with significant mass loss and suspect that this may be related to the numerical challenges in modeling the stellar envelope discussed in Paxton et al. (2013).
3.3 Angular momentum
We now discuss the limit in the parameter space which gives rise to either a final increase or increase then final diminishment in the spin angular momentum of the massive star. For simulations with an initial three day orbital period in Fig. 2, the third row gives the change in spin angular momentum (solid line) in units of stellar angular momentum at break-up velocity , where . For both metallicities, the spin-up is characterized by an increase then a sharp fall-off after the peak coincident with mass loss due to RLOF given in the second row. For simulations with an initial twenty-one day orbital period in Fig. 3, this behavior is also shown in the high metallicity () case. The magnitude of the peak is significantly smaller in the initial orbital period of twenty-one days than three days because the tidal interaction is much weaker due to the larger orbital separation.